Here is a bivariate data set.

x y
44.1 57.5
28.1 30.4
19.3 21.1
19.4 7.6
23.9 22.7
23.6 25.6
36.5 43.3
47.5 56



Find the correlation coefficient and report it accurate to three decimal places.
r =

What proportion of the variation in y can be explained by the variation in the values of x? Report answer as a percentage accurate to one decimal place.
r² = %
Answer #1
  • Here Let X be the random variable denoting the independent variable and let Y be the random variable denoting the dependent variable .
  • First we will calculate the sum of square of X and Y
  • SST=8013.94-(8*30.052)=789.92
  • y-ny=10914.9-(8*33.02)=2189.7
  • ssTy=9219.8-(8*30.05*33.0)=1280.59
  • Hence to calculate the correlation coefficient

  • SSry VSS* SSyy=

  • =0.973

  • To calculate the coefficient of determination

  • proportion of the variation in y can be explained by the variation in the values of x

  • SScy SSrr * SSyy

  • 0.948

  • =94.8%
Answer #2

Here is a bivariate data set.

xy
6229.4
77.376
39.316.8
45.231.2
50.328.9
69.865.6
79.865
64.752.2



Find the correlation coefficient and report it accurate to three decimal places. 
r = 

What percentage of the variation in y can be explained by the variation in the values of x? Report answer as a percentage accurate to one decimal place.
r² = %
Answer #3

ased on the data shown below, calculate the regression line (each value to at least two decimal places)

y =  x + 

xy
425.22
524.75
625.98
727.31
828.44
929.97
1029.5
1132.03
1234.66
1335.39
Answer #4

ased on the data shown below, calculate the regression line (each value to two decimal places)

y =  x + 

xy
325.01
426.28
527.05
630.82
732.39
832.96
938.43
1036.9
1138.87
1241.64
1345.21
1447.98
1548.65
1652.82
1754.19
1854.16